# MoC: Mixtures of Text Chunking Learners for Retrieval-Augmented Generation System

### v1

@@ A.5 Prompt utilized in Chunking
When preparing datasets using GPT-4o and generating chunking rules with MoC, prompts are necessary, as illustrated in Tables 6 and 7. The design and implementation of these prompts are crucial, as they directly influence the quality and characteristics of the resulting datasets and chunking rules.
A.6 Details of MoC Training
@@
During the model training phase, we adopt specific parameter configurations. Specifically, the training batch size per device is set to 3, and model parameters are updated every 16 steps through the gradient accumulation strategy. The learning rate is set to $51.0e-$ 5 to achieve fine-grained adjustment of weights. The model underwent a total of 3 epochs of training. Additionally, we employ a cosine annealing learning rate scheduling strategy and set a warmup ratio of 0.1 to facilitate model convergence. The variations in training loss are recorded, namely, Figure 3 showcases the training loss of the router, while Figure 5-8 individually depict the training losses of chunking experts in different intervals. The bf16 format is enabled during training to balance memory consumption and training speed. This training is conducted on two NVIDIA A800 80G graphics cards, ensuring efficient computing capabilities.
@@
[…]
@@
Figure 3: Trend of loss change during router training.

### v2

@@ A.8 Prompt utilized in Chunking
When preparing datasets using GPT-4o and generating chunking rules with MoC, prompts are necessary, as illustrated in Tables 10 and 11. The design and implementation of these prompts are crucial, as they directly influence the quality and characteristics of the resulting datasets and chunking rules.
A.9 Details of MoC Training
@@
During the model training phase, we adopt specific parameter configurations. Specifically, the training batch size per device is set to 3, and model parameters are updated every 16 steps through the gradient accumulation strategy. The learning rate is set to $51.0e-$ 5 to achieve fine-grained adjustment of weights. The model underwent a total of 3 epochs of training. Additionally, we employ a cosine annealing learning rate scheduling strategy and set a warmup ratio of 0.1 to facilitate model convergence. The variations in training loss are recorded, namely, Figure 6 showcases the training loss of the router, while Figure 8-11 individually depict the training losses of chunking experts in different intervals. The bf16 format is enabled during training to balance memory consumption and training speed. This training is conducted on two NVIDIA A800 80G graphics cards, ensuring efficient computing capabilities.
Figure 6: Trend of loss change during router training.
@@
If you understand, please segment the following text into text chunks, with each chunk enclosed using <chunk> and </chunk>. Output the complete set of segmented chunks without omissions.